Please you help me fins these problem.
How do you know?
1.) For her aquarium, Susan can choose from 4 types of fish and 3 types of plants. If she chooses one type of fish and one type of plant, how many different aquariums can Susan set up?
2.) Lottery numbers in a particular state consist of 6 digits. Each lottery ball that is drawn has six sides numbered 1-6. How many different lottery numbers are possible?
3.) The soccer team is silk-screening T-shirts. They have 4 different T-shirts and 2 different colors of ink. How many different T-shirts can be made using one ink color on a T-shirt?
4.) A travel agent is offering a vacation package. Participants choose the type of tour, a meal plan, and a hotel class from the table below

TOUR MEAL Hotel
Walking Restaurant 4-star
Boat Picnic 3-star
Bicycle 2-star
1-star
How many different vacation packages are offered?
Please rely me back as soon as you could.
Thank you so much.

Answer by stanbon(75887)

1.) For her aquarium, Susan can choose from 4 types of fish and 3 types of plants. If she chooses one type of fish and one type of plant, how many different aquariums can Susan set up?
---------------
4*3 = 12
=======================
2.) Lottery numbers in a particular state consist of 6 digits. Each lottery ball that is drawn has six sides numbered 1-6. How many different lottery numbers are possible?
6^6 = 46656
---------------------
3.) The soccer team is silk-screening T-shirts. They have 4 different T-shirts and 2 different colors of ink. How many different T-shirts can be made using one ink color on a T-shirt?
4*2= 8
====================

4.) A travel agent is offering a vacation package. Participants choose the type of tour, a meal plan, and a hotel class from the table below
TOUR MEAL Hotel
Walking Restaurant 4-star
Boat Picnic 3-star
Bicycle 2-star
1-star
How many different vacation packages are offered?
---------
Your listing is too confusing.
Ans: (# of tours)*(# of mean plans)*(# of hotel classes) = your answer
===================================
Cheers,
Stan H.